OpenStudy (anonymous):

You have a credit card with 20% apr, and an outstanding balance of $300. You haven't made any new purchases in the last three months. This particular card uses the ADB method of calculating interest. What would your finance charge be with this month if there are no hidden fees?

jimthompson5910 (jim_thompson5910):

Assuming you don't make any purchases or payments at all, your balance is going to sit at $300. Then when it comes to calculating interest, the ADB (average daily balance) is $300 which means P = 300 The interest rate is 20%, telling us that r = 20/100 = 0.20 The time in years is t = 1/12 ------------------------------------------------------- Put this all into the simple interest formula: I = P*r*t to get I = P*r*t I = 300*0.20*1/12 I = 5 which means that the credit card company earned $5 this month, which is the finance charge (another way to say interest).

OpenStudy (anonymous):

Okay, What does the P R and T stand for?

OpenStudy (anonymous):

Sorry, just got it. Could you help with 2 more please?!

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

P = principal r = rate t = time

jimthompson5910 (jim_thompson5910):

principal is the amount borrowed or loaned out rate is the interest rate (in decimal form) time is usually in years

OpenStudy (anonymous):

What about: You have a credit card with 8.9% APR that calculates interest using the ADB excluding new purchases method. The balance after your last billing cycle was $550. Whats your finance charge?

jimthompson5910 (jim_thompson5910):

same idea, but now P = 550 r = 0.089 t = 1/12

OpenStudy (anonymous):

So just use the same formula with different numbers?

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

Okay, and the last one is also a little different. You have a credit card with 22.3% APR that calculates interest using the ADB including new purchases method. Last month, you bought a stereo for $1,000 on the second day of your billing cycle. Your billing cycle is thirty-one days. Before the purchase, your balance was $200. What will your finance charge be?

jimthompson5910 (jim_thompson5910):

"Before the purchase, your balance was $200" and "you bought a stereo for $1,000 on the second day of your billing cycle"

jimthompson5910 (jim_thompson5910):

Those two pieces of info mean that you had a balance of $200 for 1 whole day The balance then jumps to 1000+200 = 1200 on day 2 The balance sits at $1200 for the remaining 30 days (since 1+30 = 31)

jimthompson5910 (jim_thompson5910):

Now let's calculate the ADB (average daily balance) (200*1 + 1200*30)/31 = 1,167.741935 = 1,167.74 So the ADB is $1,167.74

jimthompson5910 (jim_thompson5910):

this is what the credit card company will use to calculate the interest (or finance) charge

jimthompson5910 (jim_thompson5910):

making sense?

OpenStudy (anonymous):

Somewhat, I just need to read over it a few times. But so far, so good for the most part!

jimthompson5910 (jim_thompson5910):

this means P = 1,167.74 r = 0.223 t = 1/12

OpenStudy (anonymous):

For that second one, would I= 4.06285?

OpenStudy (anonymous):

Okay, I see it now.

jimthompson5910 (jim_thompson5910):

second one?

OpenStudy (anonymous):

Yes, this one " You have a credit card with 8.9% APR that calculates interest using the ADB excluding new purchases method. The balance after your last billing cycle was $550. Whats your finance charge?" And for that last one would the answer be I=21.613699? Also, are these numbers dollars or percentages or what?

jimthompson5910 (jim_thompson5910):

550*0.089*1/12 = 4.079166 = $4.08

jimthompson5910 (jim_thompson5910):

1167.74*0.223*1/12 = 21.70050 = $21.70 idk how you got your answers, so double check things

OpenStudy (anonymous):

So I'm doing this incorrectly? Or am I supposed to round?

jimthompson5910 (jim_thompson5910):

you're close, but a bit off

OpenStudy (anonymous):

Okay, as long as I'm close!!

jimthompson5910 (jim_thompson5910):

well it's always good to be as accurate as possible (at least to two decimal places)

jimthompson5910 (jim_thompson5910):

you got 21.61 when you should get 21.70

jimthompson5910 (jim_thompson5910):

so you're off by 9 cents, not a big deal, but it adds up

OpenStudy (anonymous):

I had to convert 1/12 into a decimal which I think is 0.083? I don't have a calculator right now that does the fractions for me

jimthompson5910 (jim_thompson5910):

oh i see, i'd stick to fractions as much as possible til you get to the final answer

jimthompson5910 (jim_thompson5910):

if you must use decimals, use larger and more accurate versions 1/12 = 0.83333333 the '3's repeat forever

OpenStudy (anonymous):

Yeah I just have trouble multiplying fractions is all. Oh, I guess I added an extra 0 by accident!

jimthompson5910 (jim_thompson5910):

hmm i see

jimthompson5910 (jim_thompson5910):

oh wait oops

jimthompson5910 (jim_thompson5910):

i meant to say 1/12 = 0.083333333

jimthompson5910 (jim_thompson5910):

the other decimal i posted was way too big

OpenStudy (anonymous):

Oh! yeah haha that makes more sense! I got a huge number with that other one. So, For one of these questions I had to see how much I wanted to save up a month and I chose $300.. Now it's asking "With a conservative interest rate of 3.25%, how much would you have in ten years?" Is there a formula for that?

jimthompson5910 (jim_thompson5910):

so you're putting $300 in a savings account for 10 yrs?

OpenStudy (anonymous):

Supposedly! Does that make any sense?

jimthompson5910 (jim_thompson5910):

yeah, assuming simple interest, you'd earn I = P*r*t I = 300*0.0325*10 I = 97.50 So you'd earn $97.50 in interest alone In total, you'd have 300+97.50 = $397.50 in the account (principal + interest)

OpenStudy (anonymous):

I would only have that much after 10 years?

jimthompson5910 (jim_thompson5910):

assuming a onetime deposit of $300, yes

OpenStudy (anonymous):

What do you mean? I'd be saving $300 a month for 10 years. I just want to make sure I understand correctly

jimthompson5910 (jim_thompson5910):

oh so you're putting $300 into the account each month for 10 yrs? and not one time only?

OpenStudy (anonymous):

Yes, figuratively speaking :)

jimthompson5910 (jim_thompson5910):

have you learned about annuities?

OpenStudy (anonymous):

I don't believe I have!

jimthompson5910 (jim_thompson5910):

hmm strange how they ask this question but not teach you annuities yet

OpenStudy (anonymous):

This is the very last part of this class and I took a break from it for months. So if I did, I don't remember a thing

jimthompson5910 (jim_thompson5910):

have a look at this http://www.investopedia.com/terms/f/future-value-annuity.asp tell me if it looks familiar

OpenStudy (anonymous):

Will do!!:) Give me a second

OpenStudy (anonymous):

That actually really doesn't look familiar, let me go to the chapter this assignment comes from to see if it's there..

jimthompson5910 (jim_thompson5910):

yeah most likely it'll be in your book

OpenStudy (anonymous):

I do these packets of work and to be honest they aren't the most helpful things in the world, if you couldn't already tell.

OpenStudy (anonymous):

See, the lesson that question comes from goes over the history of credit, credit abuse recovery, debt elimination plans, etc.. no annuity

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the question?

OpenStudy (anonymous):

Financial contracts, credit reports, loan basics.. Yeah I'll have to scan a picture of it so hold on

jimthompson5910 (jim_thompson5910):

ok, you can either scan it or use a camera (phone) to get the digital pic of it

OpenStudy (anonymous):

Yeah, scanning is just easier! It goes directly to my files

OpenStudy (anonymous):

Should be just another minute..

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

ok thanks

OpenStudy (anonymous):

Yup:)

jimthompson5910 (jim_thompson5910):

so yeah it's $300 per month for the next 10 years

jimthompson5910 (jim_thompson5910):

ie 12*10 = 120 months

jimthompson5910 (jim_thompson5910):

without interest, you'd put away 120*300 = $36,000 into your savings

jimthompson5910 (jim_thompson5910):

however, interest complicates this problem

jimthompson5910 (jim_thompson5910):

which is where annuities come in

OpenStudy (anonymous):

And would you understand this by chance? It's the last assignment I need to do.

OpenStudy (anonymous):

Yeah, like I said it's really not the best..

jimthompson5910 (jim_thompson5910):

I'm not sure what to make of the second pic

OpenStudy (anonymous):

That's okay, I'll just have to really ask a parent lol

jimthompson5910 (jim_thompson5910):

but the general rule is that government backed investments or savings (like US treasury bonds or your FDIC bank account) are the safest investment options. However, the return is dismal compare this to high risky things like stocks which provide a high return, but they could crash and you'd be left with nothing

jimthompson5910 (jim_thompson5910):

basically: the higher the risk, the higher the reward/return

OpenStudy (anonymous):

I'll keep that in mind. So I'd have to know annuity for the other one?

jimthompson5910 (jim_thompson5910):

yes, on that page I sent you

OpenStudy (anonymous):

Okay, I just don't know what the letters mean! What do they stand for?

jimthompson5910 (jim_thompson5910):

C = cashflow (ie the amount you put in each month) C = 300

OpenStudy (anonymous):

Found it

jimthompson5910 (jim_thompson5910):

i = interest rate per period i = 0.0325/12 = 0.00270833333333 (approximate)

jimthompson5910 (jim_thompson5910):

n = number of periods (aka number of payments) n = 10*12 = 120 so you'll deposit 120 payments of $300 (think of a deposit as a payment to your savings account, that's how I think of it)

OpenStudy (anonymous):

okay, and once i plug that in is that the answer?

jimthompson5910 (jim_thompson5910):

yep, tell me what you get

OpenStudy (anonymous):

Okay, I don't think I have a good enough calculator righ now, I'll get back to you tomorow with what I have! I have to go :( Thanks!!

jimthompson5910 (jim_thompson5910):

ok cya later. You can always use an online calculator (which is free)

OpenStudy (anonymous):

Yeah, I ended up doing that. By the way, the answer I got is $153,240.67

jimthompson5910 (jim_thompson5910):

This is what I got \[\Large FV = C * \left[ \frac{(1+i)^{n} - 1}{i} \right] \] \[\Large FV = 300 * \left[ \frac{(1+0.00270833333333)^{120} - 1}{0.00270833333333} \right] \] \[\Large FV = 42,471.4441213281 \] \[\Large FV = 42,471.44 \] So there will be $42,471.44 in the account in 10 years.